# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v1

@@ 1.2 Notation
  • Graph intersections and unions. For $H,S\subset\mathcal{K}_{n}$, denote by $H\cap S$ the graph with vertex set given by $V(H)\cap V(S)$ and edge set given by $E(H)\cap E(S)$. Denote by $S\cup H$ the graph with vertex set given by $V(H)\cup V(S)$ and edge set $E(H)\cup E(S)$. In addition, denote by $S\Cap H$, $S\mathbin{\setminus\mkern-5.0mu\setminus}H$ and S
@@
    the graph induced by the edge set $E(S)\cap E(H)$, $E(S)\setminus E(H)$ and $E(S)\triangle E(H)$, respectively (in particular, these induced graphs have no isolated points).
@@ 2 Low-degree framework and algorithmic contiguity
Initially emerged from the works of [5, 32, 31, 30], the low-degree polynomial framework was later refined and extended in various directions. It has found applications in a wide range of problems, including detection problems such as planted clique, planted dense subgraph, community detection, sparse-PCA, correlated random graphs (see [32, 31, 30, 37, 50, 12, 2, 19, 43, 15, 36]), optimization problems such as maximal independent sets in sparse random graphs [23, 53], and constraint satisfaction problems such as random $k$-SAT [6] (see also the survey [37]). Furthermore, it is conjectured in [30] that the failure of degree-$D$ polynomials implies the failure of all “robust” algorithms with running time $n^{\widetilde{O}(D)}$ (here $\widetilde{O}$ means having at most this order up to a $\operatorname{poly}\log n$ factor). In the remaining of this paper, we will focus on applying this framework in the context of high-dimensional hypothesis testing problems.
@@
Roughly speaking, the approach of low-degree polynomials amounts to studying the capabilities and limitations of algorithms that can be represented as low-degree polynomial functions of the input variables (in our case, the entries of $\mathsf{Y}$). More precisely, denote $\mathcal{P}_{D}=\mathcal{P}_{n,D}$ for the set of polynomials from $\mathbb{R}^{N}$ to $\mathbb{R}$ with degree no more than $D$. With a slight abuse of notation, we will often say “a polynomial” to mean a sequence of polynomials $f=f_{n}\in\mathcal{P}_{n,D}$, one for each problem size $n$; the degree $D=D_{n}$ of such a polynomial may scale with $n$. The polynomial degree is thought of as a measure of the algorithm’s complexity and a proxy for runtime, with degree $D$ corresponding to runtime roughly $n^{D}$ (up to $\log n$ factors in the exponent), which is the number of terms in such a polynomial. As suggested by [30], the key quantity is the low-degree advantage
@@
The low degree conjecture, proposed in [30], can be summarized as follows.
@@
This informal conjecture is inspired by [30, Hypothesis 2.1.5 and Conjecture 2.2.4], as well as the fact that low-degree polynomials capture the best known algorithms for a wide variety of statistical inference tasks. The conjecture appears to hold up for distributions $\mathbb{P},\mathbb{Q}$ of a particular style that often arises in high-dimensional statistics, and we refer the readers to [30, 37, 35, 56] for further discussion on which distributions $\mathbb{P},\mathbb{Q}$ are appropriate for this framework. In addition, we point out that although in most applications (and in the statement of [30, Hypothesis 2.1.5]) we will usually take $\mathbb{Q}$ to be a “null” measure and $\mathbb{P}$ to be a “planted” measure (which makes (2.1) more tractable), several recent works [49, 34] showed that this framework might also be applicable for many “planted-versus-planted” problems. Nevertheless, we will be more conservative and will explicitly indicate in this paper whenever $\mathbb{Q}$ is treated as a planted measure.
The framework in Conjecture 2.2 provides a useful tool for probing the computational feasibility of strong or weak detection. However, as discussed in Subsection 1.1, it turns out that the failure of strong detection is not enough in our cases, especially when we hope to perform some reductions between statistical models in a regime where weak (but not strong) detection is possible. Thus, in this regime, we aim to characterize a stronger framework that rules out all one-sided test. This motivates thus to propose the following notion of algorithmic contiguity, which is defined as follows.
@@
We say that $\mathbb{Q}$ is algorithmic contiguous with respect to $\mathbb{P}$, denoted as $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$, if no $\mathbb{Q}$-based one-sided testing algorithm runs in time $n^{\widetilde{O}(D)}$. We say that $\mathbb{Q}$ and $\mathbb{P}$ are algorithmic mutually contiguous, denoted as $\mathbb{Q}\bowtie_{\leq D}\mathbb{P}$, if both $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$ and $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$ hold.
Recall that in probability theory we say a probability measure $\mathbb{Q}=\mathbb{Q}_{n}$ is contiguous with respect to $\mathbb{P}=\mathbb{P}_{n}$, if for all sequence of events $\{A_{n}\}$ we have $\mathbb{P}(A_{n})\to 0$ implies that $\mathbb{Q}(A_{n})\to 0$. Thus, our definition can be regarded as the generalization of contiguity in algorithmic view. Our main result in this section can be stated as follows.
@@
Assuming the Conjecture 2.2, for the high-dimensional hypothesis testing problem between $\mathbb{P}$ and $\mathbb{Q}$, if $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ and $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then we have $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$.
The main part of this section is devoted to the proof of Theorem 2.4. We first briefly explain our proof ideas. Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ with
[…]
${}\mathbb{P}^{\prime}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)\mbox{ and }\mathbb{Q}^{\prime}(\mathcal{A}(\mathsf{Y})=0)=1-\epsilon\mbox{ for some }\epsilon=\epsilon_{n}\to 0\,.$ (2.4)
Define
${}M=M_{n}=\epsilon_{n}^{-\frac{1}{2}}\,.$ (2.5)
Definition 2.5.
Now assuming that (2.4) holds, we see that
[…]
¯$\overline{\mathbb{P}}\Big{(}\big{(}\mathcal{A}(\mathsf{Y}_{1}),\ldots,\mathcal{A}(\mathsf{Y}_{M})\big{)}\neq(0,\ldots,0)\Big{)}\geq\Omega(1)\,.$ (2.7)
Thus, there is an efficient algorithm that achieves weak detection between $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$. Our next result, however, shows the low-degree advantage $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}\big{)}$ is bounde by $1+o(1)$.
Lemma 2.6.
Proof.
Note that
$\frac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\overline{\mathbb{P}}(\cdot\mid\kappa=i)}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}(\mathsf{Y}_{i})\,.$ (2.8)
In addition, recall (2.1). If we identify $\mathcal{P}_{D}$ as a (finite-dimensional) Hilbert space with the inner product $\langle f,g\rangle=\mathbb{E}_{\mathbb{Q}^{\prime}}[fg]$, then using Schmidt orthogonalization procedure we can find a standard orthogonal basis $\{f_{\alpha}:\alpha\in\Lambda\}$ of this Hilbert space. In particular, we can choose $\Lambda$ such that $0\in\Lambda$ and $f_{0}=1$. Note that in some cases (e.g., when $\mathbb{Q}^{\prime}$ is a product measure), this standard orthogonal basis have a closed form; however, for general $\mathbb{Q}^{\prime}$ the explicit form of $\{f_{\alpha}\}$ is often intractable. We first show that
${}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=\Bigg{(}\sum_{\alpha\in\Lambda}\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}(\mathsf{Y})]^{2}\Bigg{)}^{\frac{1}{2}}\,.$ (2.9)
Indeed, for any $f\in\mathcal{P}_{D}$, it can be uniquely expressed as
$f=\sum_{\alpha\in\Lambda}C_{\alpha}f_{\alpha}\,,$
with equality holds if and only if $C_{\alpha}\propto\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}]$. This yields (2.9). Now, note that ¯$\overline{\mathbb{Q}}=(\mathbb{Q}^{\prime})^{\otimes M}$ is a product measure of $\mathbb{Q}^{\prime}$, there is a natural standard orthogonal basis under $\overline{\mathbb{Q}}$, given by
Thus, similarly as in (2.9), we see that
$\Big{(}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}\big{)}\Big{)}^{2}=\sum_{\begin{subarray}{c}(\alpha_{1},\ldots,\alpha_{M}):\alpha_{i}\in\Lambda\\ \sum_{i=1}^{M}\operatorname{deg}(f_{\alpha_{i}})\leq D\end{subarray}}\mathbb{E}_{\overline{\mathbb{P}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\Big{]}^{2}\,.$ (2.10)
In addition, using (2.8), we see from direct calculation that
=(2.8)${=}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}_{\overline{\mathbb{Q}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\cdot\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}(\mathsf{Y}_{i})\Big{]}$
=$=\begin{cases}1\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0)\,;\\ \frac{1}{M}\mathbb{E}_{\mathbb{P}^{\prime}}\big{[}f_{\alpha_{j}}(\mathsf{Y}_{j})\big{]}\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0,\alpha_{j},0,\ldots,0)\,;\\ 0\,,&\mbox{otherwise}\,.\end{cases}$ (2.11)
Plugging (2.11) into (2.10), we get that
[…]
where in the second equality we use (2.9) and the assumption that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$. This completes our proof. ∎
We can now finish the proof of Theorem 2.4.
Proof of Theorem 2.4.
Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ satisfying (2.3). Consider the hypothesis testing problem stated in Definition 2.5. Using (2.6) and (2.7), we see that there is an efficient algorithm that achieves weak detection between $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$, which contradicts with Lemma 2.6 and the equivalence between Items (1) and (2) in Conjecture 2.2. ∎
@@
In the remaining part of this work we will also need a strengthening of Conjecture 2.2, as introduced in Subsection 1.1. As discussed in the beginning of this section, the low-degree conjecture asserts that (for certain testing problems) low-degree polynomials are at least as powerful as all algorithms of the corresponding runtime (where the correspondence is described in Conjecture 2.2). For our purposes, we introduce a natural refinement of the low-degree conjecture: we posit that low-degree polynomials perform at least as well as all algorithms of the corresponding runtime in terms of the value of the ratio on the right-hand side of (2.1).
@@
    There is no algorithm with running time $n^{\widetilde{\Theta}(D)}$ that achieves strong detection (respectively, weak detection) between $\mathbb{P}$ and $\mathbb{Q}$.
@@
    For all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{\widetilde{\Theta}(d)}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=O(1)$ (respectively, $\mathsf{Adv}(f)=1+o(1)$), where
    $\mathsf{Adv}(f)=\frac{\mathbb{E}_{\mathbb{P}^{\prime}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}^{\prime}}[f^{2}]}}=O(1)\,.$
Note that in Conjecture 2.7 we are allowed to replace $\mathbb{P},\mathbb{Q}$ with some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ that are statistically indistinguishable with $\mathbb{P},\mathbb{Q}$. This modification enables us to avoid some situations where the low-degree advantage explodes due to some “rare events” (see [3, 15, 12, 9] for example). This revised low-degree conjecture was first proposed in [44] (in a slightly different manner) for a different purpose where they aimed to study the limits of precise error of all computationally feasible algorithms. However, we will show in the next two sections that this conjecture is also useful in performing reductions between different inference tasks.
@@ 3 Partial recovery in correlated random graphs
In this section, we will use the framework we established in Section 2 to show the hardness of partial matching in correlated random graphs, thus justifying Items (1) and (2) in Theorem 1.6. To this end, we first state the precise meaning of an algorithm achieves partial matching.
@@
We say an algorithm $\mathcal{A}$ achieves weak partial matching, if it takes $(A,B)$ as input and outputs a family of estimators $f_{1},\ldots,f_{n}$ satisfying Items (1), (2) above and
@@
We point out that in the above definition $h_{i,j}$ can be thought as the estimator of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$, where Item (3) implies this algorithm correctly matches a positive fraction of $\pi_{*}(i)$ with probability 1(1)111-(1)1 - ( 1 ) (or with positive probability) and Items (1) and (2) are some regularity requirements. Our result in this section can be stated as follows:
@@
    If $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon,s)$ and let $\mathbb{P}_{*}$ to be the joint low of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $\lambda,k,\epsilon,s$ satisfy the assumptions in Item (2) of Theorem 3.2. Then there exists an event $\mathcal{E}$ such that $\mathbb{P}_{*}(\mathcal{E})=\omega(1)$ such that for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{\widetilde{\Theta}(D)}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}\mid\mathcal{E}]=o(1)$ for all $1\leq i\leq n$.
Clearly, based on Proposition 3.3, we can deduce Theorem 3.2 via a simple Markov inequality. The rest of this section is devoted to the proof of Proposition 3.3. In the following subsections, our main focus is on proving Item (1) of Proposition 3.3. Given the similarity between the proofs of Item (1) and Item (2), for Item (2) we will provide an outline with the main differences while adapting arguments from proving Item (1) without presenting full details.
@@
We first introduce some notations used in [15].
@@
We remark here that our definition of “bad” amounts to an atypically large edge density, with a carefully chosen quantitative threshold on “large”. Roughly speaking, we expect that any subgraph with size no more than $D^{2}=n^{o(1)}$ of a sparse Erdős-Rényi graph has edge-to-vertex ratio $1+o(1)$. In the definition of $\Phi$, the term $n^{1+4/D}D^{20}$ should be thought as $n^{1+o(1)}$, and $qD^{6}$ as $n^{-1+o(1)}$. The $o(1)$ terms are tuned carefully so that for a typical subgraph $H$ of a sparse Erdős-Rényi graph, $\Phi(H)$ is much bigger than $1$. The choice of $(\log n)^{-1}$ as the $\Phi$-threshold for bad graph is somewhat arbitrary, which we will only need to be vanishing as $n\to\infty$. In [15], the authors showed that one the one hand, we have $\mathbb{P}_{*}(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mathbb{P}_{*},\overline{\mathbb{P}}_{*}),\operatorname{TV}(\mathbb{P},\overline{\mathbb{P}})=o(1)$; on the other hand, we have $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$, thus verifying the low-degree hardness for the detection problem. The first step of our proof is to slightly generalize the result in [15], as incorporated in the following lemma.
@@
The proof of Lemma 3.5 is quite technical and thus is postponed to the end of this subsection. Now, based on Lemma 3.5, we establish the following result, which basically suggests that it is impossible to obtain a “good approximation” of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$ in some sense.
@@
In addition, it is straightforward to check that for all $i\in\Lambda$
$\mathbb{E}_{\overline{\mathbb{P}}}\big{[}(1-g_{i})^{2}\mid\pi_{*}(1)=i\big{]}\leq n\cdot\mathbb{E}_{\mathbb{P}_{*}}\big{[}(\mathbf{1}_{\{\pi_{*}(1)=i\}}-g_{i})^{2}\big{]}\leq 1-\tfrac{c}{2}\,.$
@@
¯$\overline{\mathbb{P}}(g_{i}>\tfrac{c}{2})\geq\Omega(1)\,.$
Since $\operatorname{TV}(\mathbb{P},\overline{\mathbb{P}})=o(1)$, we have
${}\mathbb{P}^{\prime}(g_{i}>\tfrac{c}{2})\geq\Omega(1)\mbox{ for all }\operatorname{TV}(\mathbb{P}^{\prime},\mathbb{P})=o(1)\,.$ (3.9)
However, using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture 2.7) we have $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$. Thus we must have
@@
This yields that

### v2

@@ 1.3 Notations
  • Graph intersections and unions. For $H,S\subset\mathcal{K}_{n}$, denote by $H\cap S$ the graph with vertex set given by $V(H)\cap V(S)$ and edge set given by $E(H)\cap E(S)$. Denote by $S\cup H$ the graph with vertex set given by $V(H)\cup V(S)$ and edge set $E(H)\cup E(S)$. In addition, denote by $S\Cap H$, $S\Cup H$ and S
@@
    the graph induced by the edge set $E(S)\cap E(H)$, $E(S)\cup E(H)$ and $E(S)\triangle E(H)$, respectively (in particular, these induced graphs have no isolated points).
@@ 2 Low-degree framework and algorithmic contiguity
The low-degree polynomial framework first emerged from the works of [BHK+19, HS17, HKP+17, Hop18] and has since been refined and extended in various directions. It has found applications across a broad spectrum of problems, including detection tasks such as planted clique, planted dense subgraph, community detection, sparse PCA, and tensor PCA (see [HS17, HKP+17, Hop18, KWB22, SW22, DMW23+, BKW20, DKW+22, MW25, DDL23+, KMW24]), optimization problems like finding maximal independent sets in sparse random graphs [GJW20, Wein22], and constraint satisfaction problems such as random $k$-SAT [BH22]; see also the survey [KWB22]. Additionally, it is conjectured in [Hop18] that the failure of degree-$D$ polynomials implies the failure of all “robust” algorithms with running time $n^{\widetilde{O}(D)}$ (here $\widetilde{O}$ means having at most this order up to a $\operatorname{poly}\log n$ factor). In the remaining of this paper, we will focus on applying this framework in the context of high-dimensional hypothesis testing problems.
@@
Roughly speaking, the low-degree polynomial approach focuses on understanding the capabilities and limitations of algorithms that can be expressed as low-degree polynomial functions of the input variables (in our case, the entries of $\mathsf{Y}$). To be more precise, let $\mathcal{P}_{D}=\mathcal{P}_{n,D}$ denote the set of polynomials from $\mathbb{R}^{N}$ to $\mathbb{R}$ with degree no more than $D$. With a slight abuse of notation, we will often refer to “a polynomial” to mean a sequence of polynomials $f=f_{n}\in\mathcal{P}_{n,D}$, where each $f_{n}$ corresponds to problem size $n$; the degree $D=D_{n}$ of such a polynomial may scale with $n$. As suggested by [Hop18], the key quantity is the low-degree advantage
@@
The low degree conjecture, proposed in [Hop18], can be summarized as follows.
@@
Motivated by [Hop18, Hypothesis 2.1.5 and Conjecture 2.2.4] as well as the fact that low-degree polynomials capture the best known algorithms for a wide variety of statistical inference tasks, Conjecture 2.2 appears to hold for distributions of a specific form that frequently arises in high-dimensional statistics. For further discussion on what types of distributions are suitable for this framework, we refer readers to [Hop18, KWB22, Kun21, ZSWB22]. In addition, we point out that although in most applications (and in the statement of [Hop18, Hypothesis 2.1.5]) it is typically to take $\mathbb{Q}$ to be a “null” measure and $\mathbb{P}$ to be a “planted” measure (which makes (2.1) more tractable), several recent works [RSWY23, KVWX23] showed that this framework might also be applicable for many “planted-versus-planted” problems. Nevertheless, in this paper, we adopt a more conservative view and will explicitly indicate whenever $\mathbb{Q}$ is treated as a planted measure.
The framework in Conjecture 2.2 provides a useful tool for probing the computational feasibility of strong or weak detection. However, as discussed in Subsection 1.1, it turns out that the failure of strong detection is not enough in our cases, especially when we hope to perform some reductions between statistical models in a regime where weak (but not strong) detection is possible. Thus, in this regime, we aim to characterize a stronger framework that rules out all one-sided test. This motivates thus to propose the following notion of algorithmic contiguity.
@@
We say that $\mathbb{P}$ is degree-$D$ algorithmic contiguous with respect to $\mathbb{Q}$, denoted as $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$, if no $\mathbb{Q}$-based one-sided testing algorithm runs in time $n^{\widetilde{O}(D)}$. We say that $\mathbb{Q}$ and $\mathbb{P}$ are degree-$D$ algorithmic mutually contiguous, denoted as $\mathbb{Q}\bowtie_{\leq D}\mathbb{P}$, if both $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$ and $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$ hold.
Recall that in probability theory we say a sequence of probability measure $\mathbb{P}=\mathbb{P}_{n}$ is contiguous with respect to $\mathbb{Q}=\mathbb{Q}_{n}$, if for all sequence of events $\{A_{n}\}$ we have $\mathbb{Q}_{n}(A_{n})\to 0$ implies that $\mathbb{P}_{n}(A_{n})\to 0$. Thus, our definition can be regarded as the generalization of contiguity in algorithmic view. Our main result in this section can be stated as follows.
@@
Assuming the Conjecture 2.2, for the high-dimensional hypothesis testing problem between $\mathbb{P}$ and $\mathbb{Q}$, if $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ and $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then we have $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$.
@@
The proof of Theorem 2.4 is incorporated in Subsection B.1 of the appendix. In the remaining part of this work we will also need a strengthening of Conjecture 2.2, as introduced in Subsection 1.1. As discussed in the beginning of this section, the low-degree conjecture asserts that (for certain testing problems) low-degree polynomials are at least as powerful as all algorithms of the corresponding runtime (where the correspondence is described in Conjecture 2.2). For our purposes, we introduce a natural refinement of the low-degree conjecture, which posits that low-degree polynomials perform at least as well as all algorithms of the corresponding runtime in terms of the value of the ratio on the right-hand side of (2.1).
@@
    For all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{\widetilde{O}(D)}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=O(1)$ (respectively, $\mathsf{Adv}(f)=1+o(1)$), where
    $\mathsf{Adv}(f)=\frac{\mathbb{E}_{\mathbb{P}^{\prime}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}^{\prime}}[f^{2}]}}\,.$
Note that in Conjecture 2.5 we are allowed to replace $\mathbb{P},\mathbb{Q}$ with some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ that are statistically indistinguishable with $\mathbb{P},\mathbb{Q}$. This modification enables us to avoid some situations where the low-degree advantage explodes due to some “rare events” (see [BAH+22, DDL23+, DMW23+] for example). [^1] In fact, as pointed out by Alexander S. Wein (who learned the following example from Ansh Nagda) in personal communication with the author, such truncation appears necessary for the conjecture to hold. Consider the sparse-PCA problem where the observation $Y=\tfrac{\lambda vv^{\top}}{\|v\|^{2}}+W$ where $W$ is an $n*n$ GOE matrix and $v\in\mathbb{R}^{n}$ having coordinates i.i.d. drawn from the sparse Rademacher prior with parameter $\rho=o(1)$. In this case, when $10<\lambda<$ 1 we have $\mathbb{E}_{\mathbb{P}}[f]/\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}=O(1)$ for all low-degree polynomials $f$ but $\mathbb{E}_{\mathbb{P}}[f]/\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}=\omega(1)$ if $f$ is a thresholding function on $v_{0}^{\top}Yv_{0}$ where $v_{0}$ is a fixed sparse $\{0,1\}$-vector. However, our truncated conjecture still holds as we can take $\mathbb{P}^{\prime}$ to be the conditional measure of $\mathbb{P}$ given $v_{0}^{\top}Yv_{0}$ is small. This revised low-degree conjecture was first proposed in [MW23+] (in a slightly different manner) for a different purpose where they aimed to study the limits of precise error of all computationally feasible algorithms. However, we will show in the next two sections that this conjecture is also useful in performing reductions between different inference tasks.
@@ 3 Partial recovery in correlated random graphs
In this section, we will use the framework we established in Section 2 to show the hardness of partial matching in correlated random graphs, thus justifying Items (1) and (2) in Theorem 1.6. To this end, we first state the precise meaning that an algorithm achieves partial matching.
@@
We say an algorithm $\mathcal{A}$ achieves weak partial matching, if it takes $(A,B)$ as input and outputs a family of estimators $\{h_{i,j}:1\leq i,j\leq n\}$ satisfying Items (1), (2) above and
@@
We point out that in the above definition $h_{i,j}$ can be thought as the estimator of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$, where Item (3) implies this algorithm correctly matches a positive fraction of $\pi_{*}(i)$ with probability $1-o(1)$ (or with positive probability) and Items (1) and (2) are some regularity requirements. Our result in this section can be stated as follows:
@@
    If $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon,s)$ and let $\mathbb{P}_{*}$ to be the joint low of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $\lambda,k,\epsilon,s$ satisfy the assumptions in Item (2) of Theorem 3.2. Then there exists an event $\mathcal{E}$ such that $\mathbb{P}_{*}(\mathcal{E})=\Omega(1)$ such that for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}\mid\mathcal{E}]=o(1)$ for all $1\leq i\leq n$.
Clearly, based on Proposition 3.3, we can deduce Theorem 3.2 via a simple Markov inequality. The rest of this section is devoted to the proof of Proposition 3.3. In the following subsections, our main focus is on proving Item (1) of Proposition 3.3. Given the similarity between the proofs of Item (1) and Item (2), for Item (2) we will provide an outline with the main differences while adapting arguments from proving Item (1) without presenting full details. [^2] After the completion of an earlier version of this draft, we realized that actually we can prove Proposition 3.3 (and thus proving Theorem 3.3) assuming only a truncated version of Conjecture 2.2 (which is weaker than Conjecture 2.5). However, this alternative proof seems to be more problem dependent so we defer it to Section C of the appendix. In the main text, we focus on proving Proposition 3.3 under the stronger assumption of Conjecture 2.5.
@@
We first introduce some notations used in [DDL23+].
@@
We remark here that our definition of “bad” amounts to an atypically large edge density, with a carefully chosen quantitative threshold on “large”. Roughly speaking, we expect that any subgraph with size no more than $D^{2}=n^{o(1)}$ of a sparse Erdős-Rényi graph has edge-to-vertex ratio $1+o(1)$. In the definition of $\Phi$, the term $n^{1+4/D}D^{20}$ should be thought as $n^{1+o(1)}$, and $qD^{6}$ as $n^{-1+o(1)}$. The $o(1)$ terms are tuned carefully so that for a typical subgraph $H$ of a sparse Erdős-Rényi graph, $\Phi(H)$ is much bigger than $1$. The choice of $(\log n)^{-1}$ as the $\Phi$-threshold for bad graph is somewhat arbitrary, which we will only need to be vanishing as $n\to\infty$. In [DDL23+], the authors showed that one the one hand, we have $\mathbb{P}_{*}(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mathbb{P}_{*},\overline{\mathbb{P}}_{*}),\operatorname{TV}(\mathbb{P},\overline{\mathbb{P}})=o(1)$; on the other hand, we have $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$, thus verifying the low-degree hardness for the detection problem. The first step of our proof is to generalize the result in [DDL23+], as incorporated in the following lemma.
@@
The proof of Lemma 3.5 is quite technical and thus we postpone it to Subsection B.2 of the appendix. Now, based on Lemma 3.5, we establish the following result, which basically suggests that it is impossible to obtain a “good approximation” of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$ in some sense.
@@
Recall that $\overline{\mathbb{P}}=\mathbb{P}(\cdot\mid\mathcal{E})$ where $\mathcal{E}$ is independent to the latent matching $\pi_{*}$ and $\mathbb{P}_{*}(\mathcal{E})=1-o(1)$. Thus, it is straightforward to check that for all $i\in\Lambda$
$\mathbb{E}_{\overline{\mathbb{P}}}\big{[}(1-g_{i})^{2}\mid\pi_{*}(1)=i\big{]}\leq[1+o(1)]n\cdot\mathbb{E}_{\mathbb{P}_{*}}\big{[}(\mathbf{1}_{\{\pi_{*}(1)=i\}}-g_{i})^{2}\big{]}\leq 1-\tfrac{c}{2}\,.$
@@
¯$\overline{\mathbb{P}}(g_{i}>\tfrac{c}{2}\mid\pi_{*}(1)=i)\geq\Omega(1)\,.$ (3.9)
However, using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture 2.5) we have
¯$\overline{\mathbb{P}}(\cdot\mid\pi_{*}(1)=i)\lhd_{\leq D}\mathbb{Q}\,.$
Thus we must have $\mathbb{Q}(g_{i}>\tfrac{c}{2})\geq\Omega(1)$. This yields that
@@
This further suggests that